Factor completely: 185x^2 + 9x – 2
using Factoring Trinomials

Answer by jim_thompson5910(13794)



Jump to Answer


Looking at the expression , we can see that the first coefficient is , the second coefficient is , and the last term is .


Now multiply the first coefficient by the last term to get .


Now the question is: what two whole numbers multiply to (the previous product) and add to the second coefficient ?


To find these two numbers, we need to list all of the factors of (the previous product).


Factors of :
1,2,5,10,37,74,185,370
-1,-2,-5,-10,-37,-74,-185,-370


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to .
1*(-370) = -370
2*(-185) = -370
5*(-74) = -370
10*(-37) = -370
(-1)*(370) = -370
(-2)*(185) = -370
(-5)*(74) = -370
(-10)*(37) = -370

Now let's add up each pair of factors to see if one pair adds to the middle coefficient :


First NumberSecond NumberSum
1-3701+(-370)=-369
2-1852+(-185)=-183
5-745+(-74)=-69
10-3710+(-37)=-27
-1370-1+370=369
-2185-2+185=183
-574-5+74=69
-1037-10+37=27



From the table, we can see that there are no pairs of numbers which add to . So cannot be factored.


===============================================================



Answer:


So doesn't factor at all (over the rational numbers).


So is prime.


Jump to Top